Not in book
Find the 25th term of the arithmetic sequence with first term 7 and common difference-2.

Answer by stanbon(75887)

Find the 25th term of the arithmetic sequence with first term 7 and common difference-2.
-------
nth term formula: a(n) = a(1) + (n-1)d
a(25) = 7 + (24)*-2
=7-48
= -41
------------
Cheers,
Stan H.